OpenStudy (anonymous):

This is Geometry: Find the volume, to the nearest hundredth, of a cone with a slant height 13 feet and radius 5 feet.…

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

well; \[V_{cone}=(1/3)(base.area)(height)\]

OpenStudy (amistre64):

base area is a circle so pi 5^2 = 25pi height has to be calculated for a rt tri with leg5 and leg13

OpenStudy (bbb911):

why are people stealing my question???

OpenStudy (anonymous):

simply plug in the given values into the equation: \[V=1/3(\pi)(r^2)((h)\]

OpenStudy (anonymous):

thankyou amistre

OpenStudy (amistre64):

sqrt(196+25) = sqrt(221) if i did it right

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

well bbb i proved u right again u were too ugly so no one answered u

OpenStudy (amistre64):

13^2 = 169 ...tho right lol

OpenStudy (bbb911):

here is the muliple choice for that question

OpenStudy (bbb911):

no lmao ur just jealous so go shave urself

OpenStudy (amistre64):

169 +25 ---- 194 sqrt(194) = height then

OpenStudy (amistre64):

openstudy rule of answering #112: If a question remains unanswered for .002 secs, it becomes free for any and all participants ;)

OpenStudy (bbb911):

A. 136.14 ft cube B. 314.16 ft cube C. 340.34 ft cube D. 408.141 ft cube

OpenStudy (amistre64):

\[V_{cone}=\frac{25pi.\sqrt{194}}{3}\] if I did it right ...

OpenStudy (bbb911):

so im guessing nobody knows the answer to the question

OpenStudy (amistre64):

there was a question?

OpenStudy (bbb911):

the question that stupid lmao put was originally mine not his or hers. and i put the multiple choice to prove it was mine

OpenStudy (amistre64):

ohhh.... someone pilfered your q ... got it.

OpenStudy (amistre64):

so; base area = 25pi; height = sqrt(13^2+5^2) Vcone = [25pi * sqrt(194)]/3 = abt 364.64...

OpenStudy (bbb911):

sooooo close its either C or D ... C is 340.34 ft cube and D. is 408.41 ft cube

OpenStudy (amistre64):

check to make sure all the info is right; you got a slant of 13 and a rad of 5?

OpenStudy (amistre64):

i see my mistake then

OpenStudy (amistre64):

\[h^2 + 5^2 = 13^2\] \[h=\sqrt{(169-25)}\] \[\h = \sqrt{144}=12]

OpenStudy (amistre64):

.... h = 12 25pi*12 ------ = 25pi *4 = 100pi, abt 314.16 3

OpenStudy (bbb911):

yes and that is correct thank you sooo much

OpenStudy (amistre64):

had to fight it tooth and nail, but I got it :)

OpenStudy (bbb911):

i know you did a good good and you got urself another fan :D

OpenStudy (amistre64):

:) thnx

OpenStudy (bbb911):

ur welcome :D